OpenStudy (anonymous):

Jami must have an average of at least a 93 in Science to remain in the Honor Society. Her first three test grades of the grading period are 92, 96, and 88. What must Jami score on her next test in order to have a least a 93 average in the class? a. 88 b. 92 c. 93 d. 96

OpenStudy (anonymous):

(92+96+88+x)/4=93 What did you get for your answer?

OpenStudy (anonymous):

Idk, what it is. This is confusing.

OpenStudy (amistre64):

in order to solve for x: take the needed average and multiply it by the number of tests that will be taken; and subtract the known tests scores from it

OpenStudy (amistre64):

A(#tests) -(sum of known test grades) = needed grade

OpenStudy (anonymous):

Is it D?

OpenStudy (anonymous):

To solve the equation algebriacally you first need to get ride of the 4 you are dividing by. So we multiply both sides by 4. 92+96+88+x = 372 Now solve for x

OpenStudy (anonymous):

D is correct

OpenStudy (anonymous):

No, I got 2 answers. C & D. I'm confused which 1 it is.

OpenStudy (anonymous):

C is incorrect. But, since it's multiple choice we can substitute to check our answer. We know the average is the sum of the tests divided by the number of tests. (92+96+88+96)/4 = ?

OpenStudy (anonymous):

Oooooooooooooh never mind, I messed up on my calculator Sorry I got it lol